OpenStudy (anonymous):

What is the equation of the line in standard form that passes through the point (1, 5) and is parallel to the line y = –3x – 3?

OpenStudy (anonymous):

Well, if it is in the form \[y=mx+c\]then the gradient, m must be equal to -3 for it to be parallel to \[y = –3x – 3\]Thus m = -3. Now, substitute in the values y = 5, x = 1: \[y=-3x+c\] \[5 = -3*1+c \] \[c =8\] Thus the equation is \[y=-3x+8\]or\[y+3x=8\]or\[y+3x-8=0\]